# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ Abstract
We analyze generalizations of algorithms based on the short-path framework first proposed by Hastings [Quantum 2, 78 (2018)], which has been extended and shown by Dalzell et al. [STOC ’22] to achieve super-Grover speedups for certain binary optimization problems. We demonstrate that, under some commonly satisfied technical conditions, an appropriate generalization can achieve super-quadratic speedups not only over unstructured search but also over a classical optimization algorithm that searches for the optimum by drawing samples from the stationary distribution of a Markov Chain. We employ this framework to obtain algorithms for problems including variants of Max-Bisection, Max Independent Set, the Ising Model, and the Sherrington Kirkpatrick Model, whose runtimes are asymptotically faster than those obtainable from previous short path techniques. For random regular graphs of sufficiently high degree, our algorithm is super-quadratically faster than the best rigorously proven classical runtimes for regular graphs. Our results also shed light on the quantum nature of short path algorithms, by identifying a setting where our algorithm is super-quadratically faster than any polynomial time Gibbs sampler, unless NP = RP. We conclude the paper with a numerical analysis that guides the choice of parameters for short path algorithms and raises the possibility of super-quadratic speedups in settings that are currently beyond our theoretical analysis.
@@ 1.1 Motivation
The prospect of quantum algorithmic speedups for combinatorial optimization has been heavily studied for more than two decades [FGGS00, FGG14, Has18a, Mon18, Mon20]. This interest is partially motivated by practical considerations, since combinatorial optimization problems are ubiquitous in scientific and industrial applications, and are a major source of computational bottlenecks [AAA+24, HGL+23, DMB+23]. A second principled motivation is that there are reasons to expect such a speedup, arising from the existence of quantum algorithms such as Grover’s search algorithm [Gro96], which enjoys a quadratic quantum speedup over brute force search. Since the best classical algorithms with rigorously provable runtimes for combinatorial optimization often reduce to (possibly structured) search over a large space of possible solutions, one may expect algorithms building on Grover Search to provide speedups for these algorithms. In recent years, this intuition has been largely confirmed with the development of quantum-accelerated versions that offer quadratic speedups for backtracking [Mon18, AK17] and branch-and-bound [Mon20, CMYP22], two of the main classical meta-algorithms used to obtain provable runtime guarantees. There has been some success in obtaining sub-quadratic speedups for combinatorial optimization algorithms based on Markov Chain Monte Carlo [WA08], and dynamic programming [ABI+19].
Despite this progress, there remain challenges towards leveraging quantum algorithms for combinatorial optimization. Firstly, there are many problems for which quadratic quantum speedups over the state-of-the-art classical approaches have not been demonstrated, including cases where the best algorithm is based on dynamic programming [FGS06], MCMC, and local search [Sch02, ST13]. The second challenge is more fundamental, as recent research indicates that the realization of quadratic quantum speedups is unrealistic due to constant-factor slowdowns compared to classical hardware manifesting from slower clock speeds, the overhead of error-correction, and the limited parallelizability of quantum algorithms. Realistic estimates for the resources required to execute a quantum algorithm on a scale where it can break-even with classical computing result in quantum runtimes exceeding many days. The viability of practical realization of a polynomial speedup increases with the degree of the speedup. Recent work indicating that the outlook for realizing a quartic speedup, when considering all overheads, can be much more realistic, requiring quantum runtimes on the scale of hours instead of days [BMN+21]. It is thus of fundamental importance to investigate whether it is possible to obtain “super-quadratic” speedups for combinatorial optimization.
@@
A simpler, but very non-trivial question, is whether there are quantum algorithms for combinatorial optimization that achieve Super-Grover speedups. That is, a super-quadratic speedup over unstructured search, but not necessarily the best classical algorithm. An important first step towards rigorously obtaining positive results in this direction was the “short path” quantum algorithm from Hastings [Has18a], which was demonstrated to solve combinatorial optimization problems with runtime $\mathcal{O}^{\ast}\left(2^{(0.5-c(n))n}\right)$, where $c(n)$ is a positively valued function of $n$. Here and in the sequel, we use the notation $\mathcal{O}^{*}(2^{h(n)})$ to indicate an upper bound on the runtime of the form $\operatorname{poly}(n)2^{h(n)}$ that holds for sufficiently large $n$. A limitation of this result was that the rigorously established bounds on $c(n)$ asymptote to 00 as $n$ increases, leading to sub-polynomial improvements over Grover search. In a followup work that built on the framework of Hastings (but using a significantly modified algorithm and analysis), Dalzell et al. [DPCB23] gave an algorithm that obtains strictly Super-Grover speedups for several combinatorial optimization problems, i.e., the algorithm achieves a runtime $\mathcal{O}^{*}\left(2^{(0.5-c)n}\right)$, where $c$ is a positive parameter independent of $n$. Our work largely builds on the algorithm of Dalzell et al. [DPCB23], which we henceforth refer to an improved short path algorithm due to its connection with Hastings’ original work.
@@
The successful demonstration of Super-Grover speedups leads to natural optimism that similar techniques could be used in principle to show super-quadratic speedups over the best known classical algorithm. There remain several challenges towards such a demonstration. On one hand, the speedups shown in [DPCB23] are only larger than quadratic by very small factors. On the other hand, the best performing classical algorithms for well-studied combinatorial optimization problems (while still exponential-time) are significantly faster than unstructured search. For instance, the well-known 3-SAT problem can be solved in time $\mathcal{O}^{*}\left(2^{0.39n}\right)$, the maximum independent set of an $n$-vertex graph can be found in time $\mathcal{O}^{*}(2^{0.258n})$, and the exact ground state of the Sherrington-Kirkpatrick model can be determined in time $\mathcal{O}^{*}\left(2^{0.45n}\right)$. As a consequence, the runtimes established in [DPCB23] are in most cases slower than the best classical algorithm. An exception to this is the problem of minimizing the energy of $k$-spin models, for which there has been limited study of classical algorithms. It is important to note that the mathematical analysis of [DPCB23] does not make much effort to optimize the parameters of the algorithm, and the actual runtime is predicted to be better than the theoretical predictions. However, we provide results from a numerical simulation of the algorithm in [DPCB23] which indicate that, even when the parameters are optimized, the speedup is likely to be insufficient by itself. For Maximum Independent Set on graphs with up to 21 vertices, the best runtime scaling (using a penalty term to enforce the constraints) achieved in our experiments is around $\mathcal{O}^{*}\left(2^{0.400n}\right)$. This is slower than the best classical algorithm despite optimizing the parameters, indicating that the frameworks in [Has18a, DPCB23] must be further generalized if genuine super-quadratic speedups are the goal.
@@
A second consideration is that many combinatorial optimization problems of industrial importance are constrained, including well-studied examples such as Maximum Independent Set, Maximum/Minimum Bisection, Vertex and Set Cover, Portfolio Optimization, and Hamiltonian Cycles. The current short path algorithms can only incorporate constraints by adding penalty terms to the cost function, an approach rarely used by the state-of-the-art algorithms for combinatorial optimization. Furthermore, the runtime of the short path algorithm scales with the number of bit-strings defined on the unconstrained solution space, which can often be much larger than the number of bit-strings that satisfy all constraints in the problem formulation. As an example, consider a combinatorial optimization problem with a constraint requiring solutions to have Hamming weight $\lfloor n^{\alpha}\rfloor$ for some $10<\alpha<$ 1. The number of feasible bit-strings is $2^{(1-\alpha)n^{\alpha}\log(n)}$, which is asymptotically smaller than $2^{n}$ by a significant margin. For such constrained problems, the short path algorithm cannot even offer a super-quadratic speedup over unstructured search (if the search is restricted to the feasible region).
@@
This paper seeks to address these limitations by analyzing a generalized version of short-path algorithms that, under some technical conditions, obtains super-quadratic speedups over classical algorithms that search the space of solutions using samples from the stationary distribution of a Markov Chain. We refer to such algorithms as Markov Chain Search, and the framework can be simply described [^1] We refer the reader to Section 2 for background on Markov Chains as follows. Suppose that our aim is to minimize a real-valued cost function $H\colon\mathcal{X}\mapsto\mathbb{R}$ for a finite set $\mathcal{X}\subset\{-1,1\}^{n}$, and let $P$ be the transition matrix of a Markov Chain that mixes to a stationary distribution $\pi$ supported on $\mathcal{X}$ in $\operatorname{poly}(n)$ steps. A Markov Chain Search algorithm using $P$ simply runs the chain to draw samples from $\pi$, and keeps track of the running minimum of the samples in terms of the cost function $H$. This minimum (and the corresponding sample) serves as an estimate of the global minimum (and minimizer) of $H$ on $\mathcal{X}$. To bound the expected runtime of this algorithm, it is sufficient to bound the expected number of samples before a global minimum is encountered. Letting $\pi^{\ast}$ denote the total probability that a sample from $\pi$ is a global minimizer of $H$, it follows that drawing $\mathcal{O}\left((\pi^{\ast})^{-1}\log(\delta^{-1})\right)$ samples from $\pi$ suffices to ensure that we encounter the global minimizer with probability at least $1-\delta$. We outline this framework in Algorithm 1.
@@
Markov Chain Search is a natural extension of unstructured search, with several advantages. Firstly, it is often possible to classically sample from distributions over the feasible set that favor lower cost solutions. A common example is sampling from the Gibbs distribution corresponding to the cost function $H$, where a solution $z\in\mathcal{X}$ is sampled with probability proportional to $\exp(-\beta H(z))$ for some parameter $\beta$ (usually called the inverse temperature). Gibbs distributions are the stationary distributions of well known chains such as the Glauber Dynamics [Gla63] (also referred to as Metropolis sampling), and the mixing times of these Markov Chains have been extensively studied in Physics and Computer Science for decades. It is evident that in a Gibbs distribution with $\beta>0$, low cost solutions are more likely than in the uniform distribution. If the global minima has significantly lower cost than most of the ensemble, this can lead to algorithms that are polynomially faster than unstructured search. We note that Dalzell et al. [DPCB23] make reference to precisely this framework when examining the possibility of faster classical algorithms for $k$-spin problems. A second advantage of Markov Chain Search arises when the feasible set is asymptotically much smaller than $2^{n}$ as discussed previously. Clearly, if there is a chain $\mathcal{M}$ that mixes in polynomial time to the uniform distribution over $|\mathcal{X}|$, the Markov Chain Search algorithm with $\mathcal{M}$ is faster than unstructured search. Even if we can only prepare a (not necessarily uniform) distribution whose support is restricted to $\mathcal{X}$, this may result in a runtime that is substantially better than unstructured search. While the classical analysis of precise runtimes using Markov Chain Search is typically challenging, we later give explicit examples of settings where this separation from unstructured search is realized.
@@ 1.2 Contributions
Our primary contribution is the formulation of the generalized short-path algorithm and the demonstration that, under certain technical conditions, it obtains a super-quadratic speedup over unstructured search. In particular we obtain quantum runtimes of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$ where $T(n)$ is the (exponential in $n$) runtime of the classical Markov Chain Search, and $c(n)>0$ is a positive parameter. When $c(n)$ is bounded below by a constant $c>0$ we say we have a true super-quadratic speedup, and when $c(n)=o(1)$ we say we have an asymptotically decaying advantage over quadratic speedup (in the vein of Hastings’ results [Has18a, Has18b, Has19a].) We go on to discuss how these conditions may be established and demonstrate explicit results in two settings.
@@
    As an example of constrained optimization, we focus on optimization over strings of fixed Hamming Weight, using a Markov Chain based on random transpositions. In this case, Markov Chain search reduces simply to a brute force search over feasible strings, i.e., those that satisfy the Hamming weight constraint. Correspondingly, we identify conditions on the cost function for which a super-quadratic speedup using the short path framework can be obtained over brute-force search restricted to feasible states. As an explicit example, we consider the Maximum-Bisection problem. We note that the new runtime by itself is not particularly interesting as the number of feasible solutions to Maximum-Bisection (i.e., the number of $n$-bit binary strings of Hamming weight $n/2$) is smaller than $2^{n}$ by a polynomial factor and thus, the runtime for Markov Chain Search is not notably faster than unstructured search for this problem. However, applying the generalized framework allows for an algorithm that is completely restricted to feasible states, and the corresponding runtime cannot be obtained using the existing frameworks found in [Has18a, DPCB23].
@@
    Our next application is to Markov Chain search with non-uniform sampling, particularly to search with Gibbs distributions prepared using the Glauber Dynamics. This analysis yields conditions for super-quadratic speedups for problems such as the Maximum Independent Set problem, the Sherrington Kirkpatrick Model, and the Ising Model on graphs of bounded degree, for which Markov Chain search outperforms unstructured search. For the maximum independent set problem on regular graphs, we demonstrate that if the degree is a sufficiently high constant, Markov Chain search with the Glauber dynamics outperforms the best known theoretical classical algorithms for this problem. We also identify a number of cases where our analysis yields algorithms with a super-quadratic advantage over Markov Chain search where the advantage is asymptotically quadratic for large $n$ (in the vein of Hastings’ original results [Has18a]).
The following theorem informally summarizes the applications of the framework.
Theorem 1.1 (Applications of Generalized Short-Path Framework (informal)).
The following optimization problems exhibit a quantum runtime of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$, where $T(n)$ is the runtime of a classical algorithm based on Markov Chain Search, and $c>0$ quantifies the degree of improvement over quadratic search:
@@
    Maximum Bisection on random graphs with constant average degree,where the Markov Chain is the transposition walk and $c(n)=\Theta(1)$.
@@
    Maximum Independent Set and Ising Model on graphs of constant maximum degree, where the Markov Chain is the Glauber Dynamics at any temperature that permits polynomial mixing, and $c(n)=\Theta(1)$.
@@
    A generalization of Maximum Bisection where the smaller partition is constrained to have $k$ nodes (where $k\leq n/2$). The Markov Chain is the transposition walk and $c(n)=\Theta\left(1/\log(n/k)\right)$.
@@
    The Sherrington-Kirkpatrick Model where the Markov Chain is the Glauber Dynamics at any temperature that permits polynomial mixing, and $c(n)=\Theta\left(1/\log(n)\right)$.
The first two items above represent a constant improvement over quadratic speedup, whereas in the latter two, the speedup decays to 00 as $n$ increases. The Markov Chain Search considered is faster than brute force search on Maximum Independent Set for random regular graphs of sufficiently high degree, and on all instances of the other problems.
From a technical point of view, while the skeleton of the framework follows that of [DPCB23] quite closely, the intermediate conditions must all be uplifted to incorporate the Markov Chain used for the base classical algorithm, and significant care is needed to obtain the generalized results. These generalizations are crucial to leverage the two main advantages of Markov Chain search over unstructured search, the ability to sample from non-uniform distributions and distributions with support restricted to feasible solutions. As we will later show, these cannot be achieved by specializations of existing frameworks. The uplift of the conditions also clarifies some aspects of the role they play in the argument. We found that approaching the original results of [DPCB23] from this new perspective led to some additional insights, which may be implicit, but are not explicitly documented in other works. We note also that the algorithm of [DPCB23] follows directly as a special case of our generalized framework by considering the bit-flip walk (or the random walk on the vertices of the hypercube). Finally, the generalized analysis in this paper allows us to greatly simplify some of the statistical mechanics arguments made in [DPCB23], and instead rely directly on some standard results from the theory of Markov Chains and this may be of separate interest.
Our techniques also allow us to shed more light on a fundamental question about the viability of true super-quadratic speedups with the short path framework. The algorithms in this paper as well as the earlier frameworks, rely on preparing a quantum state whose overlap with the global minimizers is larger than that of some easily prepared starting state and jumping to the global minimum from that state. It is apparent that if there existed a classical algorithm to sample in polynomial time from a distribution with overlap that matches that of this intermediate state, then there is a classical algorithm that finds the global minimum only quadratically slower than the short path algorithm. The advantage over search is then essentially “dequantized” and there is no hope for true super-quadratic speedups. It is therefore important to understand to what degree classical sampling techniques can approach the overlap of the intermediate state, as discussed in [DPCB23]. Most natural sampling algorithms are based on the analysis of Markov Chains and so our framework provides a useful tool to probe this question. Our results on the Maximum Independent Set for bounded degree graphs yields some concrete evidence that the advantage of the short path framework over search cannot simply be removed by classical sampling. Specifically, we demonstrate a concrete optimization problem for which the intermediate state prepared by the short path algorithm has higher overlap than any Gibbs distribution that can be prepared in polynomial time, unless NP = RP.

### v2

@@ Abstract
We analyze generalizations of quantum algorithms based on the short path framework first proposed by Hastings [Quantum 2, 78 (2018)], which has been extended and shown by Dalzell et al. [STOC ’23] to achieve super-Grover speedups for certain binary optimization problems. We demonstrate that, under some commonly satisfied technical conditions, an appropriate generalization can achieve super-quadratic speedups not only over unstructured search but also over a classical optimization algorithm that searches for the optimum by drawing samples from the stationary distribution of a Markov chain. We employ this framework to obtain algorithms for problems including variants of Max Bisection, Max Independent Set, and finding the ground states of the Antiferromagnetic Ising Model and the Sherrington-Kirkpatrick Model, whose runtimes are asymptotically faster than those obtainable with previous short path techniques. In certain cases, our algorithms achieve super-quadratic speedups compared to the best known classical algorithms with rigorously established runtimes. We conclude the paper with a numerical analysis that guides the choice of parameters for short path algorithms and raises the possibility of super-quadratic speedups in settings that are currently beyond our theoretical analysis.
@@ 1.1 Motivation
The prospect of quantum algorithmic speedups for combinatorial optimization has been heavily studied for more than two decades [FGGS00, FGG14, Has18a, Mon18, Mon20]. This interest is partially motivated by practical considerations, since combinatorial optimization problems are ubiquitous in scientific and industrial applications, and are a major source of computational bottlenecks [AAA+24, HGL+23, DMB+23]. A second principled motivation is that there are reasons to expect such a speedup, arising from the existence of quantum algorithms such as Grover’s search algorithm [Gro96], which enjoys a quadratic quantum speedup over brute force search. Since the best classical algorithms with rigorously provable runtimes for combinatorial optimization often reduce to (possibly structured) search over a large space of possible solutions, one may expect algorithms building on Grover search to provide speedups for these algorithms. In recent years, this intuition has been largely confirmed with the development of quantum accelerated versions that offer quadratic speedups for backtracking [Mon18, AK17] and branch-and-bound [Mon20, CMYP22], two of the main classical meta-algorithms used to obtain provable runtime guarantees. There has been some success in obtaining sub-quadratic speedups for combinatorial optimization algorithms based on Markov chain Monte Carlo (MCMC) [WA08], and dynamic programming [ABI+19] methods.
Despite this progress, there remain challenges towards leveraging quantum algorithms for combinatorial optimization. Firstly, there are many problems for which quadratic quantum speedups over the state-of-the-art classical approaches have not been demonstrated, including cases where the best algorithm is based on dynamic programming [FGS06], MCMC, and local search [Sch02, ST13]. The second challenge is more fundamental, as recent research has identified challenges towards the realization of quadratic quantum speedups due to constant-factor slowdowns compared to classical hardware manifesting from slower clock speeds, the overhead of error correction, and the limited parallelizability of quantum algorithms. Realistic estimates for the resources required to execute a quantum algorithm on a scale where it can break-even with classical computing can result in quantum runtimes exceeding many days. The viability of practical realization of a polynomial speedup increases with the degree of the speedup. For example, the resource analysis in [BMN+21] indicates that the outlook for realizing a quartic speedup, when considering all overheads, can be much more realistic, requiring quantum runtimes on the scale of hours instead of days. It is thus of fundamental importance to investigate whether it is possible to obtain “super-quadratic” speedups for combinatorial optimization.
@@
A simpler, but very nontrivial question, is whether there are quantum algorithms for combinatorial optimization that achieve super-Grover speedups. That is, a super-quadratic speedup over unstructured search, but not necessarily the best classical algorithm. An important first step towards rigorously obtaining positive results in this direction was the quantum short path algorithm from Hastings [Has18a], which was demonstrated to solve combinatorial optimization problems with runtime $\mathcal{O}^{\ast}\left(2^{(0.5-c(n))n}\right)$ [^1] We use the notation $\mathcal{O}^{*}(2^{h(n)})$ to indicate an upper bound on the runtime of the form $\mathcal{O}(\operatorname{poly}(n)2^{h(n)})$. , where $c(n)$ is a positively valued function of $n$. A limitation of this result was that the rigorously established bounds on $c(n)$ asymptote to 0 as $n$ increases, leading to sub-polynomial improvements over Grover search. In a followup work that built on the framework of Hastings (but using a significantly modified algorithm and analysis), Dalzell et al. [DPCB23] gave an algorithm that obtains strictly super-Grover speedups for several combinatorial optimization problems, i.e., the algorithm achieves a runtime $\mathcal{O}^{*}\left(2^{(0.5-c)n}\right)$, where $c$ is a positive parameter independent of $n$. Our work largely builds on the algorithm of Dalzell et al. [DPCB23], which we henceforth refer to as an improved short path algorithm due to its connection with Hastings’ original work.
@@
The successful demonstration of super-Grover speedups leads to natural optimism that similar techniques could be used in principle to show super-quadratic speedups over the best known classical algorithm. There remain several challenges towards such a demonstration. On one hand, the speedups shown in [DPCB23] are only larger than quadratic by very small factors. On the other hand, the best performing classical algorithms for well-studied combinatorial optimization problems (while still exponential-time) are significantly faster than unstructured search. For instance, even when restricting ourselves to algorithms with provable runtimes, the well-known 3-SAT problem can be solved in time $\mathcal{O}^{*}\left(1.32065^{n}\right)$ [HMS10, PPSZ05], the maximum independent set of an $n$-vertex graph can be found in time $\mathcal{O}^{*}(1.1996^{n})$ [XN17], and the exact ground state of the Sherrington-Kirkpatrick model can be determined in time $\mathcal{O}^{*}\left(1.3670^{n}\right)$ [Mon20]. As a consequence, the runtimes established in [DPCB23] are in most cases slower than the best classical algorithm. An exception to this is the problem of minimizing the energy of $k$-spin models, for which there has been limited study of classical algorithms. It is important to note that the mathematical analysis of [DPCB23] does not make much effort to optimize the parameters of the algorithm, and the actual runtime is predicted to be better than the theoretical predictions. However, we provide results from a numerical simulation of the algorithm in [DPCB23] which indicate that, even when the parameters are optimized, the speedup is likely to be insufficient by itself. For maximum independent set (MIS) problem on graphs with up to 21 vertices, the best runtime scaling (using a penalty term to enforce the constraints) achieved in our experiments is around $\mathcal{O}^{*}\left(1.3195^{n}\right)$. This is slower than the best classical algorithm despite optimizing the parameters, indicating that the frameworks in [Has18a, DPCB23] must be further generalized if genuine super-quadratic speedups are the goal.
@@
A second consideration is that many combinatorial optimization problems of industrial importance are constrained, including well-studied examples such as finding maximum independent set, maximum/minimum bisection, vertex and set cover, portfolio optimization, and Hamiltonian cycles. The current short path algorithms can only incorporate constraints by adding penalty terms to the cost function, an approach rarely used by the state-of-the-art algorithms for combinatorial optimization. Furthermore, the runtime of the short path algorithm scales with the number of bit-strings defined on the unconstrained solution space, which can often be much larger than the number of bit-strings that satisfy all constraints in the problem formulation. As an example, consider a combinatorial optimization problem with a constraint requiring solutions to have Hamming weight $\lfloor n^{\alpha}\rfloor$ for some $0<\alpha<1$. The number of feasible bit-strings is about $2^{(1-\alpha)n^{\alpha}\log(n)}$, which is asymptotically smaller than $2^{n}$ by a significant margin. For such constrained problems, the short path algorithm cannot even offer a super-quadratic speedup over unstructured search (if the search is restricted to the feasible region).
@@
This paper seeks to address these limitations by analyzing a generalized version of short path algorithms that, under some technical conditions, obtain super-quadratic speedups over classical algorithms that search the space of solutions using samples from the stationary distribution of a Markov chain. We refer to such algorithms as Markov chain search (MCS), and the framework can be simply described [^2] We refer the reader to Section 2 for background on Markov chains. as follows. Suppose that our aim is to minimize a real-valued cost function $H\colon\mathcal{X}\rightarrow\mathbb{R}$ for a finite set $\mathcal{X}\subset\{-1,1\}^{n}$, and let $P$ be the transition matrix of a Markov chain that mixes to a stationary distribution $\pi$ supported on $\mathcal{X}$ in $\operatorname{poly}(n)$ steps. A Markov chain search algorithm using $P$ simply runs the chain to draw samples from $\pi$, and keeps track of the running minimum of the samples in terms of the cost function $H$. This minimum (and the corresponding sample) serves as an estimate of the global minimum (and minimizer) of $H$ on $\mathcal{X}$. To bound the expected runtime of this algorithm, it is sufficient to bound the expected number of samples before a global minimum is encountered. Letting $\pi^{\ast}$ denote the total probability that a sample from $\pi$ is a global minimizer of $H$, it follows that drawing $\mathcal{O}\left((\pi^{\ast})^{-1}\log(\epsilon^{-1})\right)$ samples from $\pi$ suffices to ensure that we encounter the global minimizer with probability at least $1-\epsilon$. We outline this framework in Algorithm 1.
@@
Markov chain search is a natural extension of unstructured search, with several advantages. Firstly, it is often possible to classically sample from distributions over the feasible set that favor lower cost solutions. A common example is sampling from the Gibbs distribution corresponding to the cost function $H$, where a solution $z\in\mathcal{X}$ is sampled with probability proportional to $\exp(-\beta H(z))$ for some parameter $\beta$ (usually called the inverse temperature). Gibbs distributions are the stationary distributions of well known chains such as the Glauber dynamics [Gla63] (also referred to as Metropolis sampling), and the mixing times of these Markov chains have been extensively studied in physics and computer science for decades. It is evident that in a Gibbs distribution with $\beta>0$, low cost solutions are more likely than in the uniform distribution. If the global minima has significantly lower cost than most of the ensemble, this can lead to algorithms that are polynomially faster than unstructured search. We note that Dalzell et al. [DPCB23] make reference to precisely this framework when examining the possibility of faster classical algorithms for $k$-spin problems. A second advantage of Markov chain search arises when the feasible set is asymptotically much smaller than $2^{n}$ as discussed previously. Clearly, if there is a chain $\mathcal{M}$ that mixes in polynomial time to the uniform distribution over $|\mathcal{X}|$, the Markov chain search algorithm with $\mathcal{M}$ is faster than unstructured search over the unconstrained space. Even if we can only prepare a (not necessarily uniform) distribution whose support is restricted to $\mathcal{X}$, this may result in a runtime that is substantially better than unstructured search. While the classical analysis of precise runtimes using Markov chain search is typically challenging, we later give explicit examples of settings where this separation from unstructured search is realized.
In both our generalized framework and [DPCB23], the advantage over quadratic comes from producing an intermediate quantum state, commonly known as a guiding state, that has an improved overlap with the optimal solutions than the initial state does. The initial state will have a measurement distribution, in the computational basis, that can be efficiently replicated classically. This intermediate state then acts as the starting point for Grover’s algorithm [Gro96] or more specifically quantum minimum finding [DH99]. Hence, if a classical algorithm can also replicate the improved overlap attained by the intermediate state, then the overall quantum speedup would only be quadratic. Hence, we refer to this “partial dequantization” as a Groverization of the short-path algorithm. It is currently unclear whether or not the applications of previous short-path frameworks can be Groverized.
@@ 1.2 Contributions
Our primary contribution is the formulation of the generalized short path algorithm and the demonstration that, under certain technical conditions, it obtains a super-quadratic speedup over Markov chain search. In particular we obtain quantum runtimes of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$ where $T(n)$ is the (exponential in $n$) runtime of the classical Markov chain search, and $c(n)>0$ is a positive parameter. When $c(n)$ is bounded below by a constant $c>0$ we say we have a true super-quadratic speedup, and when $c(n)>0$ for $n<\infty$ but $c(n)=o(1)$ we say we have an asymptotically decaying advantage over quadratic speedup (in the vein of Hastings’ results [Has18a, Has18b, Has19]). We go on to discuss how these conditions may be established and demonstrate explicit results. We can further decompose the implications of our main result into two major contributions.
Contribution 1 – Quantum Improvements over Non-trivial Classical Algorithms :
The chosen applications highlight two major advantages of our framework. These properties enable speedups over non-trivial classical algorithms and new quantum runtimes, both of which would not be possible with the existing short path frameworks [Has18a, DPCB23].
@@
    Handling Hard Constraints: As a first example of constrained optimization, we focus on optimization over strings of fixed Hamming weight, a problem we call MaxCut Hamming, using a Markov chain based on random transpositions. In this case, Markov chain search reduces simply to a brute force search over feasible strings, i.e., those that satisfy the Hamming weight constraint. We note that the new runtime by itself is not particularly interesting as the number of feasible solutions (i.e., the number of $n$-bit binary strings of Hamming weight $n/2$) is smaller than $2^{n}$ by a polynomial factor and thus, the runtime for Markov chain search is not notably faster than unstructured search for this problem. The second constrained problem we consider is finding the ground state of the hardcore model, i.e. Maximum Independent Set, where we utilize a constraint-preserving Glauber dynamics chain. For random regular graphs of sufficiently high degree, the Glauber chain asymptotically outperforms brute-force random search over all $2^{n}$ bit-strings [^3] However, due to known computational barriers in Gibbs sampling, unlike the Maxcut Hamming case, we are unable to sample from the uniform distribution over feasible strings..
@@
    Non-uniform Prior: Our next application is to Markov chain search for unconstrained problems but with non-uniform stationary distributions, particularly to search with Gibbs distributions prepared using the Glauber dynamics [^4] While the hardcore model is a constrained problem that also falls into this category, the inability to sample from positive inverse-temperatures due to computational barriers implies that the advantage MCS has over bruteforce search comes from restricting to the space of independent sets. Thus, it is less about the non-uniformity of the prior.. The problems we consider in this setting are the antiferromagnetic Ising model and the Sherrington-Kirkpatrick model. In both cases, the non-uniform Gibbs sampler outperforms algorithms that start from a uniform distribution.
The following theorem informally summarizes our results for the above applications.
Theorem 1.1 (Applications of Generalized Short Path Framework (informal Theorem 3.5)).
The the following optimization problems exhibit a quantum runtime of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$, where $T(n)$ is the runtime of a classical algorithm based on Markov chain search, and $c(n)>0$ quantifies the degree of improvement over quadratic speedup:
@@
    solving a generalization of MaxBisection, which we call MaxCut-Hamming, over Erdős-Rényi random graphs with constant average degree,
@@
    finding the ground state of the hardcore model, i.e. maximum independent set, on graphs with constant maximum degree,
@@
    finding the ground state of the antiferromagnetic Ising model on graphs of constant maximum degree,
@@
    and finding the ground state of the Sherrington-Kirkpatrick model.
Additionally, we provide some partial evidence that it may be possible to eventually show that some of the super-quadratic speedups attained in this work cannot be Groverized.
Contribution 2 – Technical Improvements with Generalized Short-Path Framework :
From a technical point of view, while the skeleton of the framework follows that of [DPCB23] quite closely, the intermediate conditions must all be uplifted to incorporate the Markov chain used for the base classical algorithm, and significant care is needed to obtain the generalized results. These generalizations are crucial to leverage the two main advantages of Markov chain search over brute force search, the ability to sample from non-uniform distributions and distributions with support restricted to feasible solutions. As we will later show, these cannot be achieved by specializations of existing frameworks. The uplift of the conditions also clarifies some aspects of the role they play in the argument. We found that approaching the original results of [DPCB23] from this new perspective led to some additional insights, which may be implicit, but are not explicitly documented in other works. We note also that the algorithm of [DPCB23] follows directly as a special case of our generalized framework by considering the bit-flip walk (or the lazy, random walk on the vertices of the hypercube). Finally, the generalized analysis in this paper allows us to greatly simplify some of the statistical mechanics arguments made in [DPCB23], and instead rely directly on some standard results from the theory of Markov chains and this may be of separate interest.